OpenStudy (anonymous):

Janine walks at an average speed of 1 mile in 15 minutes , what is janines average speed in miles per hour? Show your work

jhonyy9 (jhonyy9):

so you know thet one hour equal 60 minutes then write 1 mile -------- 15 minutes x mile ------ 60 minutes --------------------------- x = 60*1/15 = 60/15 = 4 x=4 so this mean thet Janine walks 4 miles per hour hope so much that is understandably sure right easy